# A Compartmental Model for Epidemiology with Human Behavior and Stochastic Effects

### v2

@@ 1. Introduction
Our model includes spread of noncompliance via mass action, and is similar to that analyzed in [6] in the ODE setting and [7, 29] in the partial differential equation (PDE) setting. We discuss the specifics of these models shortly. Pant et al. [27] propose a similar model, where populations are divided into subclasses based on adherence with NPIs, but the transfer between the classes is modeled using linear terms, as opposed to mass action terms. We opt for the latter, though it complicates the analysis, because of the mechanistic precedent from social contagion theory. Other approaches for including behavioral effects in mathematical epidemiology include modeling of heterogeneous risk aversion (wherein there is behavioral diffusion) [5], kinetic models including imperfect adherence with NPIs [11], behavioral changes based on local disease incidence [10], and age-stratified models whose implicit assumption is the populations of different ages will exhibit significantly different behaviors [4, 24, 26]. For a recent reflection on human behavior in epidemiological models, see [14].
@@
Parallel to the work incorporating human behavior, there has been research on the stochastic compartmental epidemic models. Earlier work by Tornatore et al. [31], Gray et al. [12], and Ji et al. [17] appended basic susceptible-infected-recovered (SIR) type models with uncertainty in the disease transmission rate. Later work on stochastic epidemic models includes analysis and/or numerical study of stochastic PDE models under a variety of different modeling decisions [1, 21, 23], and study of seasonality of disease emergence in stochastic ODE models [16]. For overviews on more general stochastic modeling of epidemics, see [2, 3]. Notably, Pang and Pardoux [25] formulate a stochastic PDE model which is age-stratified, hence implicitly incorporating some heterogeneous behavior. However, to the best of the authors’ knowledge, this manuscript is the first to incorporate behavioral effects into a compartmental model in such a mechanistic manner, while also considering stochastic perturbation. Our model could be seen as an extension of those in [17, 31], but the addition of behavioral effects significantly complicates the analysis due to the presence of competing nonlinear terms.
@@
The remainder of this manuscript is organized as follows. In section 2, we present a deterministic ODE model for epidemiology which includes NPIs and noncompliant behavior, and give a full analysis of the stability of disease free equilibrium points for the deterministic model. In section 3, we append the model with stochastic perturbations, and prove that solutions of our new system exist globally and remain positive as long as there is positive initial data. In section 4, we provide stability analysis for the disease free equilibria of the stochastic system, and draw analogies between the deterministic and stochastic settings. Our main strategy for establishing asymptotic behavior of the stochastic system will be the construction of suitable stochastic Lyapunov functions in different parameter regimes. In section 5, we demonstrate all of our results with simulations, and we finish with brief concluding remarks and comments of avenues of future work in section 6.
@@ 2. A deterministic SIR model with noncompliant behavior
Following the modeling strategy of [6, 7, 29, 30], we imagine a scenario where the government enacts non-pharmaceutical intervention (NPI) measures, such as mask-wearing, social distancing, or shelter-at-home orders, which decrease the infection rate by decreasing the amount of mixing by some portion $\alpha\in[0,1]$. Hence, the mass action term $SI$ becomes $(1-\alpha)S\cdot(1-\alpha)I$ since individuals in both classes are mixing at a decreased rate. However, we also assume there is a portion of the population which will not comply with these measures and thus do not receive the reduction in infectivity. Thus we split each compartment $S,I,R$ into compliant individuals (who retain the labels $S,I,R$) and noncompliant individuals (henceforth labeled $S^{*},I^{*},R^{*}$). The noncompliant individuals do not decrease their level of mixing, and thus do not receive the $(1-\alpha)$ multiplier when they participate in the mass action terms corresponding to disease spread. Borrowing from social contagion theory, we imagine that noncompliance also spreads like a disease running parallel to the actual disease. We let noncompliance also spread via mass action: when complaint individuals come in contact with noncompliant individuals, they become noncompliant with some rate $\mu>0$. Social scientists also posit that beliefs and opinions naturally die out or regress to mean over time, so we allow noncompliant individuals to “recover” and become compliant again at constant rate $\nu\geq 0.$ We further assume that any newly added members of the population are noncompliant with probability $\xi\in[0,1]$. Defining the actively mixing infectious population by $I^{(M)}=(1-\alpha)I+I^{*}$ and the total noncompliant population $N^{*}=S^{*}+I^{*}+R^{*}$, All of these modeling decisions lead to the equations
@@
(2.6) $\dot{x}=\mathcal{F}(x)-\mathcal{V}(x).$
where $\mathcal{F},\mathcal{V}:{\mathbb{R}}^{6}\to{\mathbb{R}}^{6}$. Here $\mathcal{F}_{i}$ contains all terms which introduce new infections into compartment $i$, and $\mathcal{V}_{i}$ contains all births and all other transfer into and out of compartment $i$. We then further decompose $\mathcal{V}(x)=\mathcal{V}^{-}(x)-\mathcal{V}^{+}(x)$, where $\mathcal{V}^{-}_{i}$ corresponds to flow out of compartment $i$ and $\mathcal{V}^{+}_{i}$ corresponds to any flow into compartment $i$ which does not introduce new infections. For our particular model, the infection function is
@@
With nonnegative initial conditions, the solutions of (2.2) will remain nonnegative, so our state space is ¯+$\overline{\mathbb{R}}^{6}_{+}$. We define the disease free (DF) set
@@
Definition 2.1 (Stability of $T\subset U_{DF}$).
We say that $T\subset U_{DF}$ is disease-free-stable (DF-stable) if there is a neighborhood $Z$ of $T$ in ¯+$\overline{\mathbb{R}}^{6}_{+}$ such that any solution $x(t)$ of (2.6) with $x(0)\in Z$ satisfies
@@
Intuitively, $T\subset U_{DF}$ is disease-free-stable if dynamics beginning near enough to $T$ tend toward $U_{DF}$ exponentially fast as $t\to\infty$. In particular, for an equilibrium point $x_{0}\in U_{DF}$, the classical notion of local asymptotic stability implies this notion disease free stability for $\{x_{0}\}$.
Given this definition, to prove DF-stability of $U_{0}$, one can linearize around a point in $U_{0}$, and analyze the flow into and out of the infectious compartments for the linearized system, as demonstrated in [20]. We take a point $x_{s,s^{*}}=(0,0,s,0,s^{*},0)\in X_{0}$ so that $s+s^{*}\leq b/\delta$. Linearizing (2.2), around this point, we find
(2.10) $D\mathcal{F}(x_{s,s^{*}})=\begin{bmatrix}\beta(1-\alpha)^{2}s&\beta(1-\alpha)s&\cdots\\ \beta(1-\alpha)s^{*}&\beta s^{*}&\\ \vdots&&\ddots\end{bmatrix}$
@@
(2.15) $\mathscr{R}_{0}(s,s^{*})=\frac{\beta}{\gamma+\delta}\left(s\left[(1-\alpha)^{2}+\frac{\alpha(1-\alpha)\mu s^{*}}{\gamma+\delta+\nu+\mu s^{*}}\right]+s^{*}\left[1-\frac{\alpha\nu}{\gamma+\delta+\nu+\mu s^{*}}\right]\right)$
This formula looks complicated, but has a fairly straightforward interpretation. The term $\frac{\beta}{\gamma+\delta}$ in front is the ratio of the disease spread and disease clearance rates: here $\beta$ is the infection rate and $\gamma+\delta$ is the clearance of the disease through either recovery or natural death. What remains is weighted by the portion of the population which is either compliant or noncompliant. Among the compliant population (represented by the term with $s$ in front), the $(1-\alpha)^{2}$ reflects a the reduction in infectivity gained by both the susceptible and infectious populations reducing their mixing, and the term $\frac{\alpha(1-\alpha)\mu s^{*}}{\gamma+\delta+\nu+\mu s^{*}}$ captures the increase in infectivity due to aggregate effects of (1) the susceptible compliant population’s interactions with the infections noncompliant population and (2) the flow of the compliant population toward noncompliance. Among the noncompliant population (represented by the term with $s^{*}$ in front), the $1$—in contrast to $(1-\alpha)^{2}$—represents the lack of reduction in infectivity in interactions between the noncompliant susceptible class and the noncompliant infectious class, and the term $\frac{\alpha\nu}{\gamma+\delta+\nu+\mu s^{*}}$ being subtracted from the $1$ represents the aggregate reduction in infectivity among the noncompliant population due to (1) the compliant infectious class reducing their mixing and (2) the recovery from noncompliance which is occurring. Notice in particular that
(2.16) $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\frac{b}{\delta}\frac{\beta(1-\alpha)^{2}}{\gamma+\delta},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}\left(1-\frac{\alpha\nu}{\gamma+\delta+\nu+\mu s^{*}}\right).$
The first is the reproductive ratio from the basic SIR model with infection rate $\beta(1-\alpha)^{2}$ when the total population is at its steady state of $b/\delta$, which makes sense because if the initial state is $(S,S^{*})=(\tfrac{b}{\delta},0)$, there is no possibility of noncompliance developing. The second encapsulates the “worst case scenario” where the population is at its steady state and the entire population is noncompliant. In this case, the reproductive ratio is essentially $\beta/(\gamma+\delta)$ (multiplied by the population), but decreased slightly due to the possibility of individuals becoming compliant.
Given this definition of $\mathscr{R}_{0}(s,s^{*})$, we would like to cite [20, 33] in order to conclude that $\mathscr{R}_{0}(s,s^{*})<1$ ensures DF-stability of $x_{s,s^{*}}$. However, because there is a possibility of the population becoming noncompliant which increases the infectivity, small perturbations from $x_{s,s^{*}}$ could lead to outbreaks of the disease if enough infectious people are introduced and the population becomes noncompliant fast enough. Mathematically, there is an assumption in [33] which in our context would require $D\mathcal{V}(x_{s,s^{*}})$ to have eigenvalues with positive real part. We address this momentarily. As a preliminary result, we can make a claim regarding stability if the largest possible $\mathscr{R}_{0}$ is small enough.
@@
This follows from lemma 2.2 and [20, Theorem 3.1] after noting that the family of linearizations (2.13) satisfy $\rho(F_{s,s^{*}}V_{s,s^{*}}^{-1})=\mathscr{R}_{0}(s,s^{*})\leq\mathscr{R}_{0}(0,\tfrac{b}{\delta})<1$, meaning that (2.13) exhibits exponential decay with decay rate which is uniform over $x_{0}\in U_{0}$. ∎
@@
Our theorem follows from [33, Theorem 2]. We use the definitions of $\mathcal{F},\mathcal{V}^{\pm}$ in (2.7),(2.8), and let $\mathcal{F}_{i}$ denote the $i^{\text{th}}$ coordinate of $\mathcal{F}$ and similarly for $\mathcal{V}^{\pm}$. With this notation, assumptions (A1)-(A5) from [33, Theorem 2] can be written:
@@
Note that conditions (A1)-(A4) must hold generally. The last condition (A5) is local to the particular DFE one is analyzing. Under these conditions, [33, Theorem 2], gives local asymptotic stability of the DFE when $\mathscr{R}_{0}<1$, and instability when $\mathscr{R}_{0}>1.$
@@
By zooming on the first principle $2\times 2$ submatrix in (2.22), we observe that there are two eigenvalues of $[D\mathcal{V}(x_{s,s^{*}})]^{T}$ which have eigenvectors of the form $(z_{1},z_{2},0,0,0,0)$. These must also be eigenvalues of $D\mathcal{V}(x_{s,s^{*}})$ and can be easily found to be
@@
$\frac{\delta+\nu}{\mu}-\sqrt{\left(\frac{b}{\delta}-\frac{\delta+\nu}{\mu}\right)^{2}+\frac{4\xi b}{\mu}}<\frac{\delta+\nu}{\mu}.$

### v4

@@ 1. Introduction
Our model includes spread of noncompliance via mass action, and is similar to that analyzed in [8] in the ODE setting and [9, 48] in the partial differential equation (PDE) setting. We discuss the specifics of these models shortly. Pant et al. [46] propose a similar model, where populations are divided into subclasses based on adherence with NPIs, but the transfer between the classes is modeled using linear terms, as opposed to mass action terms. We opt for the latter, though it complicates the analysis, because of the mechanistic precedent from social contagion theory. Other approaches for including behavioral effects in mathematical epidemiology include modeling of heterogeneous risk aversion (wherein there is behavioral diffusion) [6], kinetic models including imperfect adherence with NPIs [20], behavioral changes based on local disease incidence [16], and age-stratified models whose implicit assumption is the populations of different ages will exhibit significantly different behaviors [5, 43, 45]. We are specifically inspired to noncompliance with NPIs as a social contagion due to reports such as [4, 34] wherein roughly 10% of individuals not adhering to NPIs list their primary reason for not adhering as some variation of ”none of my friends are adhering,” suggests that peer imitation and social pressure play a significant role in NPI adoption. For a recent reflection on incorporation of human behavior in epidemiological models, see [25].
@@
Parallel to the work incorporating human behavior, there has been research on the stochastic compartmental epidemic models. Earlier work by Tornatore et al. [51], Gray et al. [22], and Ji et al. [28] appended basic susceptible-infected-recovered (SIR) type models with uncertainty in the disease transmission rate. Later work on stochastic epidemic models includes analysis and/or numerical study of stochastic PDE models under a variety of different modeling decisions [1, 36, 42], and study of seasonality of disease emergence in stochastic ODE models [27]. For overviews on more general stochastic modeling of epidemics, see [2, 3]. Notably, Pang and Pardoux [44] formulate a stochastic PDE model which is age-stratified, hence implicitly incorporating some heterogeneous behavior.
This work lies at the intersection of compartmental models for epidemiology incorporating human behavior à la [8] and those incorporating stochastic effects à la [22, 28, 51]. To the best of the authors’ knowledge, this manuscript is the first to model behavioral effects in such a mechanistic manner, while also considering stochastic perturbation. As such, this is an extension of the work in each of [8, 22, 28, 51].
@@
The remainder of this manuscript is organized as follows. In section 2, we present a deterministic ODE model for epidemiology which includes NPIs and noncompliant behavior, and give a full analysis of the stability of disease free equilibrium points for the deterministic model. In section 3, we append the model with stochastic perturbations, and prove that solutions of our new system exist globally and remain positive as long as there is positive initial data. In section 4, we provide stability analysis for the disease free equilibria of the stochastic system, and draw analogies between the deterministic and stochastic settings. Our main strategy for establishing asymptotic behavior of the stochastic system will be the construction of suitable stochastic Lyapunov functions in different parameter regimes. In section 5, we demonstrate all of our results with simulations, and we finish with brief concluding remarks and comments regarding avenues of future work in section 6.
As a final note before commencing, we mention that for our stochastic model in sections 3 and 4, we opt to perturb the deterministic model using Gaussian white noise with the Itô interpretation as in [22, 28, 51]. This is to distinguish from the Stratonovich calculus approach used, for example, in [7, 12, 19, 32]. Braumann [10] elucidates the difference between these approaches, and concludes that, in essence, it is “merely semantic,” stemming from the informal interpretation of phrases like average per capita growth rate or average infection rate. In the context of SIR type models, the difference usually presents itself mathematically in the following way. When using Itô noise, the strength of perturbation typically appears directly in the conditions for disease extinction or persistence as seen below and in [22, 28, 51]. By contrast, when using Stratonovich noise, the randomness does not cause a change in the reproductive ratio or the extinction criteria, but rather changes the bounds on the infected population in the case of disease persistence; see, for example [19, 32]. The Itô and Stratonovich interpretations have some theoretical tradeoffs as well. For example, in the Itô formulation for any reasonably well-behaved function $f$, the integral $\int^{t}_{0}fdW_{t}$ is a continuous local martingale, while the same is not true in the Stratonovich formulation. However, a stochastic ODE with Stratonovich noise can be easily realized as a limit of a specific type of perturbed deterministic ODE via the Wong-Zakai theorem [18, 32], while the same is not true when using Itô noise. In another direction, some authors introduce noise via an Ornstein-Uhlenbeck process [33, 38] which has the mean-reverting property making the possibility of negative perturbation less likely [39, 40], as indeed d’Onofrio [17] points out that in some biological models, Gaussian white noise can lead to questionable results. However, we choose white noise to reflect that random effects can potentially raise or lower infectivity rates, and neither our theorems nor our simulations indicate biologically absurd results. Instead, our results imply that while the random perturbations can raise or lower infectivity, to ensure disease extinction, the policy-maker must “plan for the worst”, though individual paths for the stochastic ODE may reflect better or worse outcomes than in the corresponding deterministic case. All this to say that while there are many manners of introducing stochastic perturbation, following [22, 28, 51], we use simple Gaussian white noise and leave the integration of other forms of randomness into our model as an avenue of future work.
@@ 2. A deterministic SIR model with noncompliant behavior
We modify (2.1) in several ways, and in doing so, introduce several new variables and parameters. Thus, in addition to the discussion in the ensuing paragraph, to aid the reader, we include descriptions of all variables and parameters in figure 1 and include a flow diagram for our system in figure 2.
Following the modeling strategy of [8, 9, 48, 49], we imagine a scenario where the government enacts non-pharmaceutical intervention (NPI) measures, such as mask-wearing, social distancing, or shelter-at-home orders, which decrease the infection rate by decreasing the amount of mixing by some portion $\alpha\in[0,1]$ for those who comply. We then split each compartment $S,I,R$ into compliant individuals (who retain the labels $S,I,R$) and noncompliant individuals (henceforth labeled $S^{*},I^{*},R^{*}$). The noncompliant individuals do not receive a reduction in mixing. Borrowing from social contagion theory, we imagine that noncompliance also spreads via mass action with rate $\mu>0$, parallel to the disease spread. While some research suggests the possibility that NPI fatigue may make it unlikely for noncompliant individuals to readopt compliant behavior [21], other work suggests that well-implemented public advocacy programs can increase compliance with NPIs [31, 50]. Accordingly, for full generality, we allow noncompliant individuals to become compliant again with transfer rate $\nu\geq 0.$ We note that this effect can be removed by setting $\nu=0$ if one prefers and all subsequent derivations and theorems still hold.
Finally, we assume that any newly added members of the population are noncompliant with probability $\xi\in[0,1]$. Defining the actively mixing infectious population by $I^{(M)}=(1-\alpha)I+I^{*}$ and the total noncompliant population $N^{*}=S^{*}+I^{*}+R^{*}$. All of these modeling decisions lead to the equations
@@
Quantity Description
$S$ compliant susceptible population
II compliant infectious population
$R$ compliant recovered population
$S^{*}$ noncompliant susceptible population
$I^{*}$ noncompliant infectious population
$R^{*}$ noncompliant recovered population
$N^{*}$ total noncompliant population ($N^{*}=S^{*}+I^{*}+R^{*})$
$b$ natural birth rate $(b>0)$
$\xi$ portion of newly introduced population which is noncompliant ($\xi\in[0,1]$)
$\delta$ natural death rate ($\delta>0$)
$\beta$ natural infectivity strength of the disease $(\beta>0)$
$\gamma$ natural recovery rate for the disease $(\gamma>0)$
$\alpha$ reduction in infectivity due to compliance with NPIs ($\alpha\in(0,1]$)
$\mu$ baseline infectivity strength of noncompliance ($\mu>0$)
$\nu$ maximal achievable rate of recovery from noncompliance ($\nu\geq 0)$
Figure 1. A list of variables and parameters for (2.2).
@@
Figure 2. The flow diagram for (2.2). Arrows leaving a given population denote outward flow proportional to that population with the given rate. Colored arrows connote population transfer which is modified or introduced for our model (i.e., manners in which our model substantively differs from the basic SIR model).
@@
(2.6) $\dot{x}=\mathcal{F}(x)-\mathcal{V}(x),$
where $\mathcal{F},\mathcal{V}:{\mathbb{R}}^{6}\to{\mathbb{R}}^{6}$. Here $\mathcal{F}_{i}$ contains all terms which introduce new infections into compartment $i$, and $\mathcal{V}_{i}$ contains all other transfer into and out of compartment $i$. We then further decompose $\mathcal{V}(x)=\mathcal{V}^{-}(x)-\mathcal{V}^{+}(x)$, where $\mathcal{V}^{-}_{i}$ corresponds to flow out of compartment $i$ and $\mathcal{V}^{+}_{i}$ corresponds to any flow into compartment $i$ which does not introduce new infections. For our particular model, the infection function is
@@
With nonnegative initial conditions, the solutions of (2.2) will remain nonnegative, so our state space is $\overline{{\mathbb{R}}}^{6}_{+}$. We define the disease free (DF) set
@@
Definition 2.1 (Stability of $U\subset U_{DF}$).
We say that $U\subset U_{DF}$ is disease-free-stable (DF-stable) if there is a neighborhood $Z$ of $U$ in $\overline{{\mathbb{R}}}^{6}_{+}$ such that any solution $x(t)$ of (2.6) with $x(0)\in Z$ satisfies
@@
Intuitively, $U\subset U_{DF}$ is disease-free-stable if dynamics beginning near enough to $U$ tend toward $U_{DF}$ exponentially fast as $t\to\infty$. In particular, for an equilibrium point $x_{0}\in U_{DF}$, the classical notion of local asymptotic stability implies this notion disease free stability for $\{x_{0}\}$.
To apply the definition as in [35], we linearize (2.2) around a point $x_{s,s^{*}}=(0,0,s,0,s^{*},0)\in X_{0}$ satisfying $s+s^{*}\leq b/\delta$. We see
(2.10) $D\mathcal{F}(x_{s,s^{*}})=\begin{bmatrix}\beta(1-\alpha)^{2}s&\beta(1-\alpha)s&\cdots\\ \beta(1-\alpha)s^{*}&\beta s^{*}&\\ \vdots&&\ddots\end{bmatrix}$,
@@
(2.15) $\mathscr{R}_{0}(s,s^{*})=\frac{\beta}{\gamma+\delta}\left(s\left[(1-\alpha)^{2}+\frac{\alpha(1-\alpha)\mu s^{*}}{\gamma+\delta+\nu+\mu s^{*}}\right]+s^{*}\left[1-\frac{\alpha\nu}{\gamma+\delta+\nu+\mu s^{*}}\right]\right).$
This formula has a fairly straightforward interpretation as a weighted average of the reproductive ratios for the compliant and noncompliant classes, complicated by the fact that there is some transfer between the two. Notice in particular that
(2.16) $\color[rgb]{0,0,0}\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\frac{b}{\delta}\frac{\beta(1-\alpha)^{2}}{\gamma+\delta},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}\left(1-\frac{\alpha\nu}{\gamma+\delta+\nu+\mu s^{*}}\right).$
The first is the reproductive ratio from the basic SIR model with infection rate $\beta(1-\alpha)^{2}$ which occurs when the entire population is compliant. The second encapsulates the “worst case scenario” where the population is at its steady state and the entire population is noncompliant. In this case we lose the $(1-\alpha)^{2}$ reduction in infectivity, but the reproductive ratio is still slightly decreased due to the possibility of individuals becoming compliant.
Given this definition of $\mathscr{R}_{0}(s,s^{*})$, we can establish stability properties for disease free equilibria.
@@
This follows from lemma 2.2 and [35, Theorem 3.1] after noting that the family of linearizations (2.13) satisfy $\rho(F_{s,s^{*}}V_{s,s^{*}}^{-1})=\mathscr{R}_{0}(s,s^{*})\leq\mathscr{R}_{0}(0,\tfrac{b}{\delta})<1$, meaning that (2.13) exhibits exponential decay with decay rate which is uniform over $x_{0}\in U_{0}$. ∎
@@
Our theorem follows from [53, Theorem 2]. We use the definitions of $\mathcal{F},\mathcal{V}^{\pm}$ in (2.7),(2.8), and let $\mathcal{F}_{i}$ denote the $i^{\text{th}}$ coordinate of $\mathcal{F}$ and similarly for $\mathcal{V}^{\pm}$. With this notation, assumptions (A1)-(A5) from [53, Theorem 2] can be written:
@@
Note that conditions (A1)-(A4) must hold generally. The last condition (A5) is local to the particular DFE one is analyzing. Under these conditions, [53, Theorem 2], gives local asymptotic stability of the DFE when $\mathscr{R}_{0}<1$, and instability when $\mathscr{R}_{0}>1.$
@@
By zooming on the first principle $2\times 2$ submatrix in (2.22), we observe that there are two eigenvalues of $D\mathcal{V}(x_{s,s^{*}})]^{T}$ which have eigenvectors of the form $(z_{1},z_{2},0,0,0,0)$. These must also be eigenvalues of $D\mathcal{V}(x_{s,s^{*}})$ and can be easily found to be
@@
$\frac{\delta+\nu}{\mu}-\sqrt{\left(\frac{b}{\delta}-\frac{\delta+\nu}{\mu}\right)^{2}+\frac{4\xi b}{\mu}}<\frac{\delta+\nu}{\mu},$